Let us consider two sections: section 1 contains 7 male and 10 female students, and section 2 contains 9 male and 7 female students. One section is selected at random and then one student is selected from this section. Find the probability that the selected student is male.


1
Expert's answer
2022-03-29T16:02:58-0400

"P=\frac{1}{2}*\frac{7}{17}+\frac{1}{2}*\frac{9}{16}=\frac{265}{544}\approx 0.4871."